Super Systems Ltd. manufactures a number of computer hardware products. The following variable cost standards have been set for one of their products. Particulars Standard Price/rate Direct materials $10 per pound Direct labor ? Variable manufacturing overhead $5.375 per hour Total standard variable cost to manufacture the product is $75 per product. During February, 850 direct labor hours were worked and 600 units of the product were manufactured. Each unit of product required one unit of raw material. The actual cost per unit manufactured during February was $0.30 higher than the standard cost. For the month of February, the standard cost of production for direct labor was $14,000 and for variable manufacturing overhead was $4,300. Actual variable manufacturing overhead was $4,340. Actual costs incurred for materials used were $24,000. The material quantity usage variance was an unfavorable $700. Fixed overhead per direct labor is $3.5 (assuming normal capacity of 900 labor hours). Actual fixed overhead incurred was $3,500 (use labor hours as cost driver for overhead costs). There was no stock of materials in the beginning of the February however, at the end of the month there were 100 units of materials. Question 1 - Calculate the total standard cost of materials used for actual production during February "$41,300 " "$18,300 " "$21,000 " "$26,700 "
Answer #1

Standard labor hours =Standard variable manufacturing overhead / Variable manufacturing overhead = 4300 / 5.375 = 800 hours

Standard direct labor rate per hour = 14000 / 800 = 17.5

Standard hours per unit = 800/ 600 = 1.33

Standard direct labor cost = 17.5 * 1.33 =23.28

Standard variable manufacturing cost = 1.33 *5.375 = 7.2

Standard direct material cost = 75 - 23.28 -7.2 = 44.5 i.e . 4.45 per pound

Standard cost of materials used for actual production = 4.45*10*600 = 26700

Option $ 26700